the midpoint of AB is M(9,7).One endpoint is A(3,9).Find the coordinates of endpoint B.
Answer by Alan3354(69443)

the midpoint of AB is M(9,7).One endpoint is A(3,9).Find the coordinates of endpoint B.
---------------
The difference in x & y have to equal the difference from A to M.
---
diffx = 9-3 = 6 --> x = 9+6 = 15
diffy = 7-9 = -2 --> y = 7-2 = 5
--> B @ (15,5)